A motorboat takes 3 hrs to make a trip downstream with a 6 mph current. Returning from the trip, on the same current, it took 5 hrs. How do I calculate the speed of the boat that's still in the water?

Answer by stanbon(75887)

A motorboat takes 3 hrs to make a trip downstream with a 6 mph current. Returning from the trip, on the same current, it took 5 hrs. How do I calculate the speed of the boat in still in water?
----
Let speed in still water be "b":
----
Downstream DATA:
time = 3 hrs ; rate = b+6 mph ; distance = r*t = 3(b+6) miles
------
Upstream DATA:
time = 5 hrs ; rate = b-6 mph ; distance = r*t = 5(b-6) miles
-------------------
Equation:
distance down = distance up
3(b+6) = 5(b-6)
3b + 18 = 5b - 30
2b = 48
b = 24 mph (speed of the boat in still water)
==========================
Cheers,
Stan H.
==============================